# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v2

@@ III Magnon-mediated interactions
hence for $\Psi_{\xi,k}$, ‘spin up’ (denoted $+$ below) and ‘spin down’ (denoted --) respectively label the majority and minority Fermi surfaces for each valley. The electron-magnon coupling in Eq. (II.4) similarly maps to
@@
In the remainder of this section we calculate the leading contributions to these scattering amplitudes coming from gapless magnon modes (neglecting contributions from gapped magnons). We treat different canting angle regimes separately, starting with the generic case.
III.1 Generic canting angles: $20<\theta_{0}<\pi/$ 2
Recall that for canting angles satisfying $20<\theta_{0}<\pi/$ 2, the only gapless mode corresponds to the $s-$ magnon branch, which one can project onto at small $q$ by sending $\delta m_{\xi}^{\phi}\rightarrow\delta m_{s-}$ and $\delta m_{\xi}^{\theta}\rightarrow 0$. The electron-magnon coupling in Eq. (II.4) accordingly projects via $\delta{\boldsymbol{m}}_{\xi}(q)\cdot{\boldsymbol{s}}_{\xi}(-q)\rightarrow\delta m_{s-}(q)S^{y}_{\xi}(-q)$. It is then straightforward to integrate out the $s-$ mode, yielding interactions
$\mathcal{S}_{\rm I}=-\frac{1}{2}\sum_{\xi_{1},\xi_{2}}\int_{q}\mathcal{G}_{s-}(q)S^{y}_{\xi_{1}}(q)S^{y}_{\xi_{2}}(-q).$ (36)
Comparing to Eq. (33) and expanding the magnon Green’s function $\mathcal{G}_{s-}(q)$ for small $q$ yields nontrivial scattering amplitudes
$g^{-+,-+}_{\xi_{1},\xi_{2}}(q)\sim-g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim\frac{\cos^{2}\theta_{0}}{2z_{s}\kappa}\frac{1}{\nu^{2}+c_{s}^{2}{{\boldsymbol{q}}}^{2}},$ (37)
@@ III.1 Generic canting angles: $20<\theta_{0}<\pi/$ 2
with $z_{s}=\cos^{2}\theta_{0}+\frac{V-J}{2J}$ a dimensionless constant. Importantly, these magnon-mediated interactions feature a divergence at low frequency and momentum—which may at first sight appear counter-intuitive. One might in particular expect the electron-Goldstone-mode coupling to scale linearly with ${\boldsymbol{q}}$, given that the electron energy is not perturbed by a ${\boldsymbol{q}}=0$ Goldstone mode [52]. Such scaling would in turn generate an extra ${\boldsymbol{q}}^{2}$ factor in the numerator of Eq. (37), killing the singularity at small $q$. Electron-electron interactions mediated by acoustic phonons—which typically take the form $g_{\rm ph}(q)\sim\frac{c_{\rm ph}^{2}{{\boldsymbol{q}}}^{2}}{\nu^{2}+c_{\rm ph}^{2}{{\boldsymbol{q}}}^{2}}$, where $c_{\rm ph}$ is the phonon velocity—provide a famous example conforming to this expectation. Yet the usual argument for linear-in-${\boldsymbol{q}}$ coupling only applies to diagonal scattering amplitudes [52]. In contrast, our pairing scenario involves off-diagonal (spin-flipping) scattering, and therefore the electron-magnon vertex can be nonvanishing at small ${\boldsymbol{q}}$, consistent with our explicit calculation [^1] For a complementary perspective on this subtle issue, including the discussion of interband electron-Goldstone coupling viewed from the rotating frame, see Ref. 57. In this work we organize our diagrammatic expansion for pairing instabilities in the original (laboratory) frame. To our knowledge, an analysis from the rotating-frame perspective has not yet been carried out for such pairing problems, but would be useful to develop in future work.. Similar behavior arises for inter-band electron-phonon coupling, which contains one less ${\boldsymbol{q}}$ factor compared to intra-band processes (see, e.g., Ref. 58).
Notice that the scattering amplitudes in Eq. (37)—which we exploit for our theory of superconductivity in Sec. IV—depend on neither $\xi_{1}$ nor $\xi_{2}$. That is, the Goldstone magnon mediates intervalley and intravalley processes with the same strength. We anticipate that intervalley scatterings directly affect the pairing interaction, as Cooper pairs naturally arise from electrons residing in opposite valleys. Intravalley scatterings will prove useful for analyzing the electron self-energy; see Sec. V.
III.2 Spin-canting phase transition: $\theta_{0}=0$
At the spin-canting phase transition where $\theta_{0}=0$, two degenerate gapless critical modes appear, corresponding to the $s-$ and $a-$ branches (see also Fig. 2). In the small-$q$ limit we can project onto these modes by replacing $\delta m_{\xi}^{\phi}\rightarrow\delta m_{s-}$ and $\delta m_{\xi}^{\theta}\rightarrow\tau_{\xi}\delta m_{a-}$, in turn projecting the electron-magnon coupling as $\delta{\boldsymbol{m}}_{\xi}(q)\cdot{\boldsymbol{s}}_{\xi}(-q)\rightarrow\delta m_{s-}(q)S^{y}_{\xi}(-q)+\delta m_{a-}(q)\tau_{\xi}S^{x}_{\xi}(-q)$. Integrating out the critical modes now generates interactions
$\mathcal{S}_{\rm I}=-\frac{1}{2}\sum_{\xi_{1},\xi_{2}}\int_{q}\mathcal{G}_{s-}(q)$ $[S^{y}_{\xi_{1}}(q)S^{y}_{\xi_{2}}(-q)$
$+\tau_{\xi_{1}}\tau_{\xi_{2}}S^{x}_{\xi_{1}}(q)S^{x}_{\xi_{2}}(-q)],$ (38)
where we used $\mathcal{G}_{s-}=\mathcal{G}_{a-}$ (which reflects degeneracy of the critical modes). In this case comparing to Eq. (33) and expanding the Green’s functions for small $q$ yields scattering amplitudes
$g^{-+,-+}_{\xi_{1},\xi_{2}}(q)$ $\sim\frac{1-\tau_{\xi_{1}}\tau_{\xi_{2}}}{2z_{s}\kappa}\frac{1}{\nu^{2}+c_{s}^{2}{\boldsymbol{q}}^{2}},$ (39)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)$ $\sim\frac{-1-\tau_{\xi_{1}}\tau_{\xi_{2}}}{2z_{s}\kappa}\frac{1}{\nu^{2}+c_{s}^{2}{\boldsymbol{q}}^{2}}.$ (40)
Apart from the $s-$ contributions already present in Eq. (37) (now with $\theta_{0}\rightarrow 0$), we obtain new $\tau_{\xi_{1}}\tau_{\xi_{2}}$ terms coming from the gapless $a-$ branch. Addition of the $a-$ mode enhances the intervalley $g^{-+,-+}_{K,K^{\prime}}$ and intravalley $g^{-+,+-}_{\xi,\xi}$ amplitudes relative to the generic canting angle case, but obliterates the intravalley $g^{-+,-+}_{\xi,\xi}$ and intervalley $g^{-+,+-}_{K,K^{\prime}}$ amplitudes. Vanishing of the latter set follows from symmetry: At $\theta_{0}=0$, the system preserves global U(1) spin rotation symmetry around the $z$ axis, which in our rotated basis enforces conservation of $S^{z}_{\rm tot}=\sum_{\xi}\int_{k}\overline{\Psi}^{\dagger}_{\xi,k}\tau_{\xi}\sigma^{z}\Psi_{\xi,k}$. Intravalley $g^{-+,-+}_{\xi,\xi}$ amplitudes encode processes that change $S^{z}_{\rm tot}$ by flipping two spins in a given valley; intervalley $g^{-+,+-}_{K,K^{\prime}}$ amplitudes similarly violate this conservation law.
III.3 Spin-orbit-coupling-free limit: $\theta_{0}=\pi/2$
@@
The ferromagnetic phase with vanishing spin-orbit coupling exhibits canting angle $\theta_{0}=\pi/2$ and supports a gapless, quadratically dispersing Goldstone mode. Following the procedure outlined above, in the small-$q$ limit we can project the electron-magnon coupling onto the gapless sector and integrate out the soft magnons. As detailed in Appendix D we thereby obtain scattering amplitudes
$g^{-+,-+}_{\xi_{1},\xi_{2}}(q)=0,~{}~{}~{}~{}g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim-\frac{1}{-i\gamma\nu+D{\boldsymbol{q}}^{2}}.$ (41)
@@
Vanishing of $g^{-+,-+}_{\xi_{1},\xi_{2}}$ descends from SU(2) symmetry present in the microscopic Hamiltonian in the $\lambda_{I}=0$ limit. In the ferromagnetic phase considered here, this symmetry is reduced to U(1), corresponding to global spin rotations about the arbitrarily selected spin polarization axis—taken here to be ${\bf\hat{x}}$. Interactions must therefore preserve the total spin along $x$, which in the rotated basis corresponds to $S^{x}_{\rm tot}=\sum_{\xi}\int_{k}\overline{\Psi}^{\dagger}_{\xi,k}\sigma^{z}\Psi_{\xi,k}$. The $g^{-+,-+}$ scattering processes flip two spins in the rotated basis in clear violation of this conservation law. The $g^{-+,+-}$ processes, by contrast, preserve the residual U(1) symmetry. Nevertheless, the latter amplitudes exhibit a gentler low-frequency divergence compared to either Eqs. (37), (39) or (40) due to the quadratic Goldstone mode dispersion emerging in the absence of spin-orbit coupling. We will see in the next section that this distinction underlies spin-orbit-enabled superconductivity in our proposed scenario.

### v3

@@ III Magnon-mediated interactions
hence for $\Psi_{\xi,k}$, ‘spin up’ (denoted $+$ below) and ‘spin down’ (denoted --) respectively label the majority and minority Fermi surfaces for each valley. The electron-magnon coupling in Eq. (II.4) maps to
@@
In the remainder of this section we calculate the leading contributions to these scattering amplitudes from the soft magnon modes $s-$ and $a-$. As preparation for the pairing problem analyzed in Sec. IV, we focus specifically on inter-valley scattering amplitudes where two electrons in the majority Fermi surface are scattered to the minority Fermi surface (see Fig. 1(c)), or vice versa—e.g., $g^{+-,+-}_{K,K^{\prime}}(q)$. Appendix D details the analysis and considers other scattering amplitudes as well. We treat different canting angle regimes separately, starting with the generic spin-orbit-coupled case where $\theta_{0}<\pi/2$.
III.1 Canting angles $\theta_{0}<\pi/2$
We obtain the magnon-mediated pairing interaction by first integrating out all magnons (see Appendix D), and then isolating the soft-magnon contributions by focusing on the leading frequency and momentum dependence. We thereby obtain intervalley scattering amplitudes
$g^{+-,+-}_{K,K^{\prime}}(q)$ $\sim\frac{1}{2z_{s}\kappa}\frac{\cos^{2}\theta_{0}}{\nu^{2}+c_{s}^{2}{\boldsymbol{q}}^{2}}+\frac{1}{2z_{a}\kappa}\frac{\cos^{2}\theta_{0}}{\nu^{2}+c_{a}^{2}{\boldsymbol{q}}^{2}+\Omega_{a}^{2}}.$ (38)
The first term on the right-hand side arises from the $s-$ Goldstone mode, which remains gapless throughout the ordered phase, while the second term arises from the low-lying $a-$ mode that becomes gapless at criticality. As a reminder, the $a-$ magnon gap is smaller than $\epsilon_{F}$ by an $\mathcal{O}(J/V)$ factor; see text below Eq. (30). As a result, pair scattering mediated by this mode features significant retardation, and therefore will be roughly as relevant for pairing as the Goldstone mode. At the critical point where $\Omega_{a}\rightarrow 0$, the restored U(1) symmetry enforces a degeneracy of the two gapless modes, which manifests as $c_{s}=c_{a}$ and $z_{s}=z_{a}$.
@@ III.1 Canting angles $\theta_{0}<\pi/2$
Intriguingly, the Goldstone-mediated part of the interaction in Eq. (38) features a divergence at low frequency and momentum—which may appear counter-intuitive. One might expect the electron-Goldstone-mode coupling to scale linearly with ${\boldsymbol{q}}$, given that the electron energy can not be perturbed by a ${\boldsymbol{q}}=0$ Goldstone mode [58]. Such scaling would in turn generate an extra ${\boldsymbol{q}}^{2}$ factor in the numerator of Eq. (38), killing the singularity at small $q$, similarly to electron-electron interactions mediated by acoustic phonons which typically take the form $g_{\rm ph}(q)\sim\frac{c_{\rm ph}^{2}{{\boldsymbol{q}}}^{2}}{\nu^{2}+c_{\rm ph}^{2}{{\boldsymbol{q}}}^{2}}$, with $c_{\rm ph}$ the phonon velocity. The resolution is that the general argument for linear-in-${\boldsymbol{q}}$ coupling only applies to diagonal scattering amplitudes [58]. In contrast, our pairing scenario involves off-diagonal (spin-flipping) scattering, and therefore the electron-magnon vertex can be nonvanishing at small ${\boldsymbol{q}}$, consistent with our explicit calculation. Similar behavior arises for inter-band electron-phonon coupling, which contains one less ${\boldsymbol{q}}$ factor compared to intra-band processes (see, e.g., Ref. 60).
Importantly, however, pairing in our scenario does not benefit from the divergent small-${\boldsymbol{q}}$ part of the pairing interaction, due to the momentum-space distance between the majority and minority Fermi surfaces. An intra-band, small-${\boldsymbol{q}}$ component can be obtained through virtual excitations from one Fermi surface to another (and back), but such effects also vanish as ${\boldsymbol{q}}\rightarrow 0$, in line with Adler’s theorem showing that long-wavelength Goldstone modes couple weakly to electrons. This subtle issue will be clarified in a forthcoming work [61].
III.2 Spin-orbit-coupling-free limit: $\theta_{0}=\pi/2$
@@
The ferromagnetic phase with vanishing spin-orbit coupling exhibits canting angle $\theta_{0}=\pi/2$ and supports a gapless, quadratically dispersing Goldstone mode. Following the procedure outlined above—which is equivalent to simply setting $\theta_{0}=\pi/2$ in Eq. (38)—we obtain
$g^{+-,+-}_{K,K^{\prime}}(q)=0.$ (39)
@@
Vanishing of $g^{+-,+-}_{K,K^{\prime}}$ descends from SU(2) symmetry present in the microscopic Hamiltonian in the $\lambda_{I}=0$ limit. In the ferromagnetic phase considered here, this symmetry is reduced to U(1), corresponding to global spin rotations about the arbitrarily selected spin polarization axis—taken here to be ${\bf\hat{x}}$. Interactions must therefore preserve the total spin along $x$, which in the rotated basis corresponds to $S^{x}_{\rm tot}=\sum_{\xi}\int_{k}\overline{\Psi}^{\dagger}_{\xi,k}\sigma^{z}\Psi_{\xi,k}$. The $g^{+-,+-}_{K,K^{\prime}}$ scattering processes flip two spins in the rotated basis in clear violation of this conservation law. This symmetry constraint prevents single-magnon exchange from mediating pairing in SU(2) ferromagnets, and provides a promising explanation for spin-orbit-proximitized superconductivity in rhombohedral graphene.